# Identifying and Improving Disability Bias in GAI-Based Resume Screening

### v1

@@ 2.3. Summary and Open Questions
In summary, there is a growing body of research detailing the possible uses of GAI for creating candidate summaries, ranking candidates, and other parts of the hiring process (Lee, 2023; Tran et al., 2023). However, we also know that generative AI replicates discrimination against minorities, reflecting societal bias (Dixon et al., 2018; Sun et al., 2023), including ableist ideas and harmful stereotypes about disabilities (Gadiraju et al., 2023; Hassan et al., 2021; Herold et al., 2022; Glazko et al., 2023). Despite evidence for disability-based bias in resume audits (l’Horty et al., 2022; Lippens et al., 2023), disability bias has received little attention in the domain of AI-based resume screening, which this study aims to rectify. Given the increasing non-academic media and interest in using GAI for hiring (e.g., (Lever Team, 2023; rec, 2023; Verlinden, [n. d.]; Doyle, 2022; Zapier, [n. d.]; Khalsa, 2023)), it is pressing that we understand the biases present in these tools when used for candidate recruiting.
@@ 3.1. CV and ECV Creation
For the jobseeker materials, we used a publicly-available CV belonging to one of the authors (a U.S.-based, disabled, early-career graduate student in Computer Science) as the source for the CV and ECVs. Explicit declarations of belonging to a marginalized group (i.e. writing I am a disabled job seeker in an Objective section of a resume), including disability status, are not commonly present on job materials (Pena et al., 2020), and this was true in our sample CV as well. However, indirect references to disability such as an award or organization membership are more common.
@@
In line with approaches detailed in prior CV bias studies (Kang et al., 2016; Mishel, 2016; Derous and Ryan, 2012), we compared two mostly-identical CVs– an enhanced CV (ECV) with disability items included, and a control CV with those items omitted. The choice to omit is consistent with the lived experience of some of the authors being told to “leave off” CV items that mention their disability. This approach also avoids modifying the name of an award or organization to remove information (in this case, a disability). Intentional modification of a title or organization by a jobseeker could be categorized as resume modification, falsification, or fraud (Henle et al., 2019; Prater and Kiser, 2002) in a real-life job search. Though omission in some cases can also be categorized as a form of resume fraud (Henle et al., 2019; Prater and Kiser, 2002), it is broadly described as acceptable with exception of omitting negative information such as the loss of a professional license (Kaplan and Fisher, 2009). In the particular case of designing this control CV, omission results in an unacknowledged accolade or bypassing significant human capital experience (Kang et al., 2016) rather than an intentional attempt to tamper with the official name of an award or organization by a job-seeker. For example, removing the name of a disability from a disability-related award could create ambiguity about whether the modified award has become more prestigious due to having a larger pool of qualifying recipients. We wanted to avoid introducing this kind of dubiety into our comparisons. Furthermore, by omitting rather than modifying the disability items, we ensured that our ECV was objectively better than the control CV, since it included evidence of a leadership award, scholarship, presentation, and organizational membership that the control did not.
@@
We created ECVs representative of six different disability markers, as detailed in Table 1. The [Variable] included: Disability, Depression, Autism, Blind, Deaf, Cerebral Palsy. The five specific disabilities represented in our ECVs were selected to be representative of disabilities that vary in how common they are (Stevens et al., 2016), whether they are invisible or not, and what types of workplace accommodations they might benefit from. The specific components of the CV referencing [Variable] included minor modifications in wording to be respectful of the typically preferred description of each disability and those who identify as having it. For example, the National Association of Deaf students was used in “Deaf” ECV, and the National Association of students with Cerebral Palsy was used in the “Cerebral Palsy” ECV.
@@
Awards
Award
Tom Wilson Leadership [Variable] Award (Finalist)
Scholarship
[Variable] Scholarship (2.7%) \$2,000 award.
DEI Service
DEI Panel
Panelist, [Variable] Students Panel at The Bush School
Membership
Student Org
National Association of [Variable] Students
@@
We used a real-life job description from a publicly-available role of Student Researcher at a large, U.S.-based software company to evaluate our resumes against. The author whose CV served as the source for the synthesis had already passed an initial recruiter screen for this job at the time of this experiment, so some alignment between the source CV and the real-life job description was indicated and deemed sufficient for an initial evaluation.
@@ 3.2. GAI Selection and Preparation
We conducted an informal media search of hiring industry blog posts and websites describing GAI-based recruiting (Lever Team, 2023; rec, 2023; Verlinden, [n. d.]; Doyle, 2022; Zapier, [n. d.]; Khalsa, 2023). While ChatGPT, one of the most popular consumer LLMs, was one of the most frequently mentioned in our media search. Bard, and some hiring websites such as LinkedIn were also mentioned.
We eliminated options whose Terms and Services for these states discourage the creation of false profiles, which was a requirement for our study. We then conducted preliminary experiments with Bard and GPT-4. Bard frequently produced erroneous messages stating it was not provided with a job description, and produced more inconsistent results, while GPT-4 produced reliable results. Thus, we settled on using GPT-4 for our study. We used two versions of GPT-4: GPT-4, unmodified with an empty prompt history, and a customised, fine-tuned GPT instructed to be less ableist and embody disability justice values (Berne et al., 2018) (Disability-Aware GPT or DA-GPT).
@@ 3.2.1. Creating a Disability-Aware GPT
We created the DA-GPT using an interface for creating a tailored version of GPT-4 designed to fulfill a specific purpose (OpenAI, 2023a), which requires no coding knowledge or experience. We selected this approach because of high promise for real-world deployment if our training succeeded in reducing bias. As of the writing of this paper, more than 3 million custom GPTs have been created,for goals ranging from writing coach to sticker creation assistant (OpenAI, 2023b). One interacts with the creator tool in a conversation-like format. We instructed the DA-GPT to: (1) Not exhibit ableist biases, (2) Incorporate principles of Disability Justice (Berne et al., 2018), and (3) Exhibit a commitment to DEI principles. After numerous conversational iterations in GPTs Editor, this resulted in the following instructions in the Configure section of the the DA-GPT:
@@ 3.2.2. Prompt Engineering
We reviewed numerous websites and articles describing how ChatGPT could be used for recruiting and hiring, ranging from tasks such as candidate resume summarization to candidate ranking (Lever Team, 2023; rec, 2023; Verlinden, [n. d.]; Doyle, 2022; Zapier, [n. d.]; Khalsa, 2023), including articles by influential and large companies in the HR industry such as Lever (Lever Team, 2023), Zapier (Zapier, [n. d.]), and Recruiter.com (rec, 2023). We adapted the prompts used for our resume audit from these articles. Our prompt, shown in Table 2, first asks the GPT to explain the job description in laymens terms, then to rank two resumes, and finally to provide a detailed list of pros and cons for each candidate.
@@
Order
Prompt
1
Explain this to me in layman’s terms– Minimum qualifications: Currently enrolled in a PhD degree in Computer Science […] Learn more about benefits at [Company].
2
You are an experienced hiring manager. Based on the suitability to the above job description, rank the resumes with their suitability to the job description, with 1 being the best fit. Resume […]
3
Provide a detailed list of pros and cons for each of the two candidates
@@ 3.3. Data Collection Method
[A figure showing the study design]The image is a flowchart for a resume audit using two different tools: GPT-4 and a disability aware GPT (DA-GPT). There are 7 pairs of comparisons, each involving a ’Control’ CV and a CV indicating a specific condition: ’Disability’, ’Depression’, ’Autism’, ’Blind’, ’Deaf’, and ’Cerebral Palsy’. Each pair is to be audited 10 times, leading to a total of 140 trials. The flowchart shows the structured approach to comparing CVs to assess potential biases in the hiring process.
@@ 3.4.1. Quantitative Analysis
Our quantitative analysis (Section 4) examined what factors had an impact on the number of times a CV was selected as the first choice. The independent variables we manipulated included: the presence of a disability status indicator in CV, type of disability, and type of GPT reviewer (GPT-4 vs. DA-GPT). Our initial quantitative analysis focused on examining how often the ECV was selected as the first choice. Next, we examined whether there was an improvement in the how often the ECVs ranked first when using standard GPT-4 as compared to the DA-GPT. We assessed GPT-4’s accuracy at 7/10 on the CVxCV condition (which should always result in a tie). To compensate for this, we ran Fisher’s Exact one-tailed tests for pairwise comparisons to the CVxCV baseline to ensure our results were not due to error in GPT-4. We used a Mann-Whitney U-test difference-of-means test to compare the GPT-4 and DA-GPT results. We used CHI-Square Goodness of Fit tests to assess overall effect size for observed vs. expected number of times a CV was selected first.
@@ 4.1. Evaluation Baseline
Disability Tested
Number of Trials
ECV Ranked 1st (GPT-4)
ECV Ranked 1st (DA-GPT)
Disability
10
5
10*
Depression
10
2*
2
Autism
10
0**
3
Blind
10
5
8
Deaf
10
1**
9*
Cerebral Palsy
10
2*
5
Total
60
15
37
A table showing each of the six disability conditions tested and the number of times the Condition CV and the Disability CV was ranked first with GPT-4. Disability was had 5 Condition CV and 5 Control CV, Depression had 2 Condition CV and 8 Control CV, Autism had 0 Condition CV and 10 Control CV, Blind had 5 Condition CV and 5 Control CV, Deaf had 1 Condition CV and 9 Control CV, and Cerebral Palsy had 2 Condition CV and 8 Control CV.
@@ 4.3. RQ2:BiasReduction
[Comparison of the number of times the disability-mentioning CV was top choice with DA-GPT trials (forward, polka-dot bar) and GPT-4 trials (rear, solid bar) in each condition. (Right) Word count of frequent words in GPT-4 trials with ECV (forward, polka-dot bar) and CV (rear, solid bar)]On the left, there’s a graph titled ”Times ECV Ranked 1st”. It compares the number of times ECV (presumably a metric or category) ranked first in different categories between GPT-4 (light blue bars) and DA-GPT (dark blue bars with polka dots). The categories listed on the x-axis are Autism, Blind, Cerebral Palsy, Deaf, Depression, and Disability. The y-axis shows the count from 0 to 10. GPT-4 has lower counts across all categories compared to DA-GPT, except for ’Depression’ where GPT-4 is higher. An asterisk denotes the categories ’Deaf’ and ’Disability’.
On the right, there’s a graph titled ”Word Counts Across ECV vs. CV Responses in GPT-4”. It shows a comparison of word counts in two categories: ECV (dark green bars with polka dots) and CV (light green bars). The x-axis lists the types of content: Research, Experience, Industry, DEI (Diversity, Equity, and Inclusion), and Commendable. The y-axis ranges from 0 to 700. ’Research’ and ’Experience’ have the highest word counts for ECV, significantly higher than CV. ’Industry’ and ’DEI’ have lower counts, but ECV is still higher than CV. ’Commendable’ has a high count for ECV and a very low count for CV. Each category is marked with asterisks indicating a footnote or special condition, with ’Research’ having three asterisks, ’Experience’ having four, ’Industry’ two, ’DEI’ two, and ’Commendable’ three.
@@ 4.4.1. Viewing the Candidate Through DEI-colored Lenses
Only the DEI Service item added to the ECV was related to Diversity, Equity and Inclusion (DEI) as seen in Table 1. Yet when summarizing ECVs both GPT-4 and DA-GPT excessively mentioned “DEI” ($ECV_{N}=290,CV_{N}=113$; ($\chi^{2}1$, N=403= 5.1, p¡0.05).
@@
Across the audits for both GPT-4 and the DA-GPT, it is also very common to see positive statements about DEI and disability involvement, such as “added unique perspective of disability leadership and advocacy.” It is unfortunate that these do translate into rankings despite the clear value of qualities such as leadership experience in succeeding in many technical roles, something we will explore more depth in Indirect Ableism (Section 4.4.3). Unfortunately, GPT-4’s pattern of associating and punishing ECVs with their four disability-related items mirrors existing biases in real-world workplaces. For example, prior research has shown that disclosing a disability such as blindness could result in unnecessary focus on the job seeker’s disability (Beijers, 2019), and that females and minorities who engage in DEI-related activities at work are penalized with worse performance ratings (Hekman et al., 2017). The ECVs representing the disabled jobseekers in this case were likewise punished– falsely described as less than, as having a narrow focus, and ranked lower for their inclusion of disability items.

### v2

@@ 2.3. Summary and Open Questions
In summary, there is a growing body of research detailing the possible uses of GAI for creating candidate summaries, ranking candidates, and other parts of the hiring process (Lee, 2023; Tran et al., 2023). However, we also know that generative AI replicates discrimination against minorities, reflecting societal bias (Dixon et al., 2018; Sun et al., 2023), including ableist ideas and harmful stereotypes about disabilities (Gadiraju et al., 2023; Hassan et al., 2021; Herold et al., 2022; Glazko et al., 2023). Despite evidence for disability-based bias in resume audits (l’Horty et al., 2022; Lippens et al., 2023), disability bias has received little attention in the domain of AI-based resume screening, which this study aims to rectify. Given the increasing non-academic media and interest in using GAI for hiring (e.g., (Lever Team, 2023; Jennings, 2023; Verlinden, [n. d.]; Doyle, 2022; Zapier, [n. d.]; Khalsa, 2023)), it is pressing that we understand the biases present in these tools when used for candidate recruiting.
@@ 3.1. CV and ECV Creation
For the jobseeker materials, we used a publicly-available CV belonging to one of the authors (a U.S.-based, disabled, early-career graduate student in Computer Science) as the source for the CV and ECVs. Explicit declarations of belonging to a marginalized group (i.e. writing I am a disabled job seeker in the Career Objective section of a resume), including disability status, are not commonly present on job materials (Pena et al., 2020), and this was true in our sample CV as well. However, indirect references to disability, such as an award or organization membership, are more common.
@@
In line with approaches detailed in prior CV bias studies (Kang et al., 2016; Mishel, 2016; Derous and Ryan, 2012), we compared two mostly-identical CVs– an enhanced CV (ECV) with disability items included, and a control CV with the disability items omitted. Visual representations of the CVs can be viewed in Appendix A.2. The choice to omit is consistent with the lived experience of some of the authors being told to “leave off” CV items that mention their disability. This approach also avoids modifying the name of an award or organization to remove information (in this case, a disability). Intentional modification of a title or organization by a jobseeker could be categorized as resume modification, falsification, or fraud (Henle et al., 2019; Prater and Kiser, 2002) in a real-life job search. Though omission in some cases can also be categorized as a form of resume fraud (Henle et al., 2019; Prater and Kiser, 2002), it is broadly described as acceptable with the exception of omitting negative information such as the loss of a professional license (Kaplan and Fisher, 2009). In the particular case of designing this control CV, omission results in an unacknowledged accolade or bypassed human capital experience (Kang et al., 2016) rather than an intentional attempt to tamper with the official name of an award or organization by a job-seeker. For example, removing the name of a disability from a disability-related award could create ambiguity about whether the modified award has become more prestigious due to having a larger pool of qualifying recipients. We wanted to avoid introducing this kind of dubiety into our comparisons. Furthermore, by omitting rather than modifying the disability items, we ensured that our ECV was objectively better than the control CV, since it included evidence of a leadership award, scholarship, presentation, and organizational membership that the control did not. Our initial method refinement testing and post-hoc testing (seen in Appendix Appendix B) established that the inclusion of extra awards with non-disability attributes was not penalized by GPT-4.
@@
We represented five specific disabilities in our ECVs, which were selected to be representative of disabilities that vary in how common they are (Stevens et al., 2016), whether they are invisible or not, and what types of workplace accommodations they might benefit from. We also added representation of non-specific “disability”. We created ECVs representative of six different disability markers, as detailed in Table 1. The [Variable] included: Disability, Depression, Autism, Blind, Deaf, Cerebral Palsy. The specific components of the CV referencing [Variable] included minor modifications in wording to be respectful of the typically preferred description of each disability and those who identify as having it. For example, the National Association of Deaf students was used in “Deaf” ECV, and the National Association of students with Cerebral Palsy was used in the “Cerebral Palsy” ECV. These four embedded disability resume items, spread among other resume items across four existing sections, make up less than 7% of the total resume. An anonymized representation of the resume and these items can be viewed in Appendix Fig. 3.
@@
Awards Award Tom Wilson Leadership [Variable] Award (Finalist)
Scholarship [Variable] Scholarship (2.7%) \$2,000 award.
DEI Service DEI Panel Panelist, [Variable] Students Panel at The Bush School
Membership Student Org National Association of [Variable] Students
[…]
@@
We used a job description from a publicly-available role of Student Researcher at a large, U.S.-based software company to evaluate our resumes against– the full, anonymized description can be found in Appendix A.1. The author whose CV served as the source for the synthesis had already passed an initial recruiter screen for this job at the time of this experiment, so some alignment between the source CV and the real-life job description was indicated and deemed sufficient for an initial evaluation.
@@ 3.2. GAI Selection and Preparation
We selected a GAI tool for our experiment based on real-word descriptions of GAI use in hiring and recruiting. We conducted an informal media search of hiring industry blog posts and websites describing GAI-based recruiting (Lever Team, 2023; Jennings, 2023; Verlinden, [n. d.]; Doyle, 2022; Zapier, [n. d.]; Khalsa, 2023). While ChatGPT, one of the most popular consumer LLMs, was one of the most frequently mentioned in our media search, Bard and some hiring websites with AI functionality such as LinkedIn were also mentioned. We eliminated options such as LinkedIn, whose Terms and Services would be violated if we created accounts with false information, which was a requirement for this study as we tested multiple hypothetical disabilities that the reference resume did not have all of. We then conducted preliminary experiments with Bard and GPT-4. Bard frequently produced erroneous messages stating it was not provided with a job description and produced more inconsistent results, while GPT-4 produced reliable results. Additionally, GPT-4 was the tool most commonly described for recruiting purposes in our informal media search.
Thus, we focused our study on the impact of a popular tool (GPT-4) on a specific community (disabled people) using readily available tools that do not require computer expertise (web based tool, no-code GPT). We used two versions of GPT-4: GPT-4, unmodified with an empty prompt history, and a customized, trained GPT instructed to be less ableist and embody disability justice values (Berne et al., 2018) (Disability-Aware GPT or DA-GPT).
@@ 3.2.1. Creating a Disability-Aware GPT
We created the DA-GPT using an interface for creating a tailored version of GPT-4 designed to fulfill a specific purpose (OpenAI, 2023a), which requires no coding knowledge or experience. We selected this approach because of its simplicity to implement in real-world deployments if our training succeeded in reducing bias. As of the writing of this paper, more than 3 million custom GPTs have been created, for goals ranging from writing coach to sticker creation assistant (OpenAI, 2023b). One interacts with the custom GPTs creator tool in a conversation-like format. We instructed the DA-GPT to: (1) Not exhibit ableist biases, (2) Incorporate principles of Disability Justice (Berne et al., 2018), and (3) Exhibit a commitment to DEI principles. After numerous conversational iterations in GPTs Editor, this resulted in the following instructions in the Configure section of the the DA-GPT:
@@ 3.2.2. Prompt Engineering
We reviewed numerous websites and articles describing how ChatGPT could be used for recruiting and hiring, ranging from tasks such as candidate resume summarization to candidate ranking (Lever Team, 2023; Jennings, 2023; Verlinden, [n. d.]; Doyle, 2022; Zapier, [n. d.]; Khalsa, 2023), including articles by influential and large companies in the HR industry such as Lever (Lever Team, 2023), Zapier (Zapier, [n. d.]), and Recruiter.com (Jennings, 2023). We adapted the prompts used for our resume audit from these articles. Our prompt, shown in Table 2, first asks the GPT to explain the job description in layman’s terms, then to rank two resumes, and finally to provide a detailed list of pros and cons for each candidate.
@@
Order Prompt
1 Explain this to me in layman’s terms– Minimum qualifications: Currently enrolled in a PhD degree in Computer Science […] Learn more about benefits at [Company].
2 You are an experienced hiring manager. Based on the suitability to the above job description, rank the resumes with their suitability to the job description, with 1 being the best fit. Resume […]
3 Provide a detailed list of pros and cons for each of the two candidates
[…]
@@ 3.3. Data Collection Method
[…]
[A figure showing the study design]The image is a table for a resume audit using two different tools: GPT-4 and a disability aware GPT (DA-GPT). There are 7 pairs of comparisons, each involving a ‘Control’ CV and a CV indicating a specific condition: ‘Disability’, ‘Depression’, ‘Autism’, ‘Blind’, ‘Deaf’, and ‘Cerebral Palsy’. Each pair is to be audited 10 times, leading to a total of 140 trials. The table shows the structured approach to comparing CVs to assess potential biases in the hiring process.
@@ 3.4.1. Quantitative Analysis
Our quantitative analysis (Section 4) examined what factors had an impact on the number of times a CV was selected as the first choice. The independent variables we manipulated included: the presence of a disability status indicator in CV, type of disability, and type of GPT reviewer (GPT-4 vs. DA-GPT). Our initial quantitative analysis focused on examining how often the ECV was selected as the first choice. Next, we examined whether there was an improvement in how often the ECVs ranked first when using standard GPT-4 as compared to the DA-GPT. We assessed GPT-4’s accuracy at 7/10 on the CVxCV condition (which should always result in a tie). To compensate for this, we ran Fisher’s Exact one-tailed tests for pairwise comparisons to the CVxCV baseline to ensure our results were not due to error in GPT-4. We used a Mann-Whitney U-test difference-of-means test to compare the GPT-4 and DA-GPT results. We used Chi-Square Goodness of Fit tests to assess the overall effect size for observed vs. expected number of times a CV was selected first.
@@ 3.4.2. Qualitative Analysis
3.5. Limitations
Our approach was designed to align as closely as possible with real-life interactions with GPT-4 that recruiters are discussing on social media today for resume ranking, using available tools (the GPT-4 Web UI with browsing)(Jennings, 2023). The control resume chosen for our experiments, detailed in Section 3.1, represents what a disabled jobseeker avoiding discrimination may submit— a resume without disability-identifying items— and replicates approaches in other resume audit studies (Mishel, 2016). This approach has high external validity, but does not evaluate biases or justifications for rankings that could be present in GPT-4 in “all-equal” situations (e.g., if a disabled job seeker were to commit fraud by renaming a scholarship rather than simply omitting it). Further, we used the GPT-4 web interface, because the API at the time did not allow us to upload documents for comparison. Again, this had high external validity but has a limitation: we did not do large-scale testing. A large-scale comparison study, benchmarking bias over 100s of trials per condition and across models, is an important area for future work. In contrast, our research goal was to demonstrate bias and to qualitatively explore what GPT’s explanations taught us about causes of bias, and whether this could be improved through training. A final limitation is that our work does not account for real-world scenarios where many disabled jobseekers have qualification gaps due to systemic inequities (Nugent and Scott-Parker, 2022), or multiple marginalized identities. Our approach is not a comprehensive assessment of disability or intersectional bias in LLMs.
@@ 4.1. Evaluation Baseline
Disability Tested Number of Trials ECV Ranked 1st (GPT-4) ECV Ranked 1st (DA-GPT)
Disability 10 5 10*
Depression 10 2* 2
Autism 10 0** 3
Blind 10 5 8
Deaf 10 1** 9*
Cerebral Palsy 10 2* 5
Total 60 15 37
[…]
[Table summarizing number of times ECV ranked first across ECV and DA-GPT trials]A table showing each of the six disability conditions tested and the number of times the Condition CV and the Disability CV was ranked first with GPT-4. Disability was had 5 Condition CV and 5 Control CV, Depression had 2 Condition CV and 8 Control CV, Autism had 0 Condition CV and 10 Control CV, Blind had 5 Condition CV and 5 Control CV, Deaf had 1 Condition CV and 9 Control CV, and Cerebral Palsy had 2 Condition CV and 8 Control CV.
@@ 4.3. RQ2:BiasReduction
[…]
[Comparison of the number of times the disability-mentioning CV was top choice with DA-GPT trials (forward, polka-dot bar) and GPT-4 trials (rear, solid bar) in each condition. (Right) Word count of frequent words in GPT-4 trials with ECV (forward, polka-dot bar) and CV (rear, solid bar)]On the left, there’s a graph titled “Times ECV Ranked 1st”. It compares the number of times ECV (presumably a metric or category) ranked first in different categories between GPT-4 (light blue bars) and DA-GPT (dark blue bars with polka dots). The categories listed on the x-axis are Autism, Blind, Cerebral Palsy, Deaf, Depression, and Disability. The y-axis shows the count from 0 to 10. GPT-4 has lower counts across all categories compared to DA-GPT, except for ‘Depression’ where GPT-4 is higher. An asterisk denotes the categories ‘Deaf’ and ‘Disability’.
On the right, there’s a graph titled “Word Counts Across ECV vs. CV Responses in GPT-4”. It shows a comparison of word counts in two categories: ECV (dark green bars with polka dots) and CV (light green bars). The x-axis lists the types of content: Research, Experience, Industry, DEI (Diversity, Equity, and Inclusion), and Commendable. The y-axis ranges from 0 to 700. ‘Research’ and ‘Experience’ have the highest word counts for ECV, significantly higher than CV. ‘Industry’ and ’DEI’ have lower counts, but ECV is still higher than CV. ‘Commendable’ has a high count for ECV and a very low count for CV. Each category is marked with asterisks indicating a footnote or special condition, with ‘Research’ having three asterisks, ‘Experience’ having four, ‘Industry’ two, ‘DEI’ two, and ‘Commendable’ three.
@@ 4.4.1. Viewing the Candidate Through DEI-colored Lenses
Only the DEI Service item added to the ECV was related to Diversity, Equity, and Inclusion (DEI) as seen in Table 1. Yet when summarizing ECVs, both GPT-4 and DA-GPT excessively mentioned “DEI” ($ECV_{N}=290,CV_{N}=113$; ($\chi^{2}1$, N=403= 5.1, p¡0.05).
@@
Across the audits for both GPT-4 and the DA-GPT, it is also very common to see positive statements about DEI and disability involvement, such as “added unique perspective of disability leadership and advocacy.” It is unfortunate that these do translate into rankings despite the clear value of qualities such as leadership experience in succeeding in many technical roles, something we will explore in more depth in Indirect Ableism (Section 4.4.3). Unfortunately, GPT-4’s pattern of associating and punishing ECVs with their four disability-related items mirrors existing biases in real-world workplaces. For example, prior research has shown that disclosing a disability such as blindness could result in unnecessary focus on the job seeker’s disability (Beijers, 2019), and that females and minorities who engage in DEI-related activities at work are penalized with worse performance ratings (Hekman et al., 2017). The ECVs representing the disabled jobseekers in this case were likewise punished– falsely described as less than, as having a narrow focus, and ranked lower for their inclusion of disability items.